Using the digits 1,2,3,4,5, and 6 ONLY ONCE, find two 3-digit whose product is as large as possible.
Answer by RAY100(1637)

Basically, we are asked to pick two 3 digit numbers, which when multiplied together are a max for this set.
.
if we try to maximize the hundreds digit:
100*200=20,000
300*400=120,000
500*600=300,000,,this suggests that the hundreds digits are 5 &6
.
working on the 10's digit:
610*520=317,200
620*510=316,200
630*540=340,200,,,,,,max,,,lets use these
640*530=339,200
.
working on the tens digits:
631*542=342,002,,,,,,max,,,,,answer
632*541=341,912
.
Therefore using only the digits 1 to 6, and each only once,
the max product is 342,002, from 631*542
.